Scot and Vidia, married taxpayers, earn $248,000 in taxable income and $6,200 in interest from an investment in City of Tampa bonds. Using the U.S. tax rate schedule for married filing jointly, how much federal tax will they owe? What is their average tax rate? What is their effective tax rate? What is their current marginal tax rate? (Do not round intermediate calculations. Round "Average tax rate" & "Effective tax rate" to 2 decimal places)
Answer #1

ANSWER

Federal tax 48099 =28179+(248000-165000)*0.24
Average Tax Rate 19.39% =48099/248000
Effective tax rate 18.92% =48099/(248000+6200)
Marginal tax rate 24%

================

DEAR STUDENT,

IF YOU HAVE ANY QUERY PLEASE ASK ME IN THE COMMENT BOX,I AM HERE TO HELP YOU.PLEASE GIVE ME POSITIVE RATING..

****************THANK YOU****************